8) (8 points) Waves: A hurricane generates waves at Bermuda 800 km south-east of Rhode Island. Two sets of waves are generate this material is asked in relation to topics of general oceanography. Please frame help in the context of general oceanography and not another earth science
Answer #1

According to the question,

frequency = 1

Velocity = 6000/8 metere per second and 6000/12 meter/ second

Wavelength of first wave =velocity / frequency = (6000/8)/1 = 750m

Wavelength of first wave =velocity / frequency = (6000/12)/1 = 500m

Now first wavelength 750 m that is 0.75km goes in 1 second

therefore, it goes 800/0.75 = 1066.67 second = 17.77 minutes

Now second wavelength 500 m that is 0.5km goes in 1 second

therefore, it goes 800/0.5 = 1600 second = 26.67 minutes